OpenStudy (anonymous):

Create your own piecewise function and graph it on the coordinate plane. Explain whether your piecewise function is continuous or discontinuous. Send both the function and the graph with your assessment. (3 points)

OpenStudy (unklerhaukus):

do you know what piecewise means?

OpenStudy (unklerhaukus):

suspose\[f(x)\equiv\left\{\begin{array}\ f_1(x)\qquad -\infty<x<a\\f_2(x) \qquad\qquad a<x<b \\f_2(x) \qquad\qquad b<x<\infty\end{array} \right.\]

OpenStudy (unklerhaukus):